Jim, Jamie, Lucia, Liz, Teeg, Jonny, Lucas
Absent: Bryn, Jesse, Arwin
Module at Hobart full or broken, but second module had key switched off
DBBC at Ke been playing up quite a few times over the last fortnight.
Tsys at Yg not working, new hardware on the way to hopefully fix.
Ke cross over switch default setting. Needs attention.
Did you know: new clean page.
Experiment name changed (e.g.) AUA003 instead of 002, make sure cal and master schedule are the same.
Ross-ter to clean kitchen.
Handover notes issues:
A few problems with *R1780*:
Hb e-remote control didn't work and neither did Dave's backup? Was this resolved and so what was the fix?
Ke DBBC problem? Is that resolved? PCI7200 BOARD. Replaced.
Patrick could you put the problem/fix you had with Ke IF power levels? Also the pointing problem and fix? Ped lost power and tripped E-stop. Needed manual reboot.
Bryn/Tiege, there were MK5 and clock problems at Yarragadee? The fix as per the handover notes seems vague.
Ross I had the same problem with multiple dim's running. After I logged in as root (su) and tried that seemed to kill them.
Could you please describe how you noticed/diagnosed that the Mark5 disk/Mark5 was faulty at Hb? 'acs problem seems to be related to testing other module.'? Didn't allow recording. Checkmk5 gave an error, scan_check 4/8th line gave a fault message. Swapped to other bank and back then it died. Module marked as bad.
“IF unit has a fault, probably with the PIC. No noise diode control at the moment, so no Tsys. Power levels into DBBC are OK though. (Jim)'” This was fixed already unless I am mistaken? IF unit works apart from noise diode control, can't measure tsys. New pic on the way.
Secondary module not correctly inserted into mk5 (I think the key was not turned). When first module became full the the mk5 gave “ERROR m5 -900 : Can't XLRappend() recording” (which is not an alarm) until noticed at 17:00UT. I changed M5 -900 to ALARM in the alarm_actions.txt for Hb.
If people aren't using Arwins comments.py script, they should be.
/home/www/auscope/opswiki/data/pages/operations/agenda2017_03_16.txt · Last modified: 2017/03/16 02:27 by Lucas Hyland